OpenStudy (anonymous):

a=8 b=8 in a right angle what is the exact value of c then give an approximation to three decimal points

OpenStudy (anonymous):

c² = a² + b² = 8² + 8² = 64 + 64 = 128 c = root(128) c = 11.3137 c = 11.314

OpenStudy (anonymous):

\[\sqrt{8^2+8^2}=\sqrt{128}=11.3137085\] 11.313

OpenStudy (anonymous):

11.314 not 11.313